OpenStudy (anonymous):

Write the slope intercept equation for the line that passes through (-8, -11 ) and is perpendicular to -2x - 7y = -14 Please show how you got the answer. Thank you!

OpenStudy (aroub):

slope intercept is y=mx+b perpendicular is the reciprocal of the slope which when u multiply them u get -1 okay till now?

OpenStudy (aroub):

okay i will continue anyways lol .... so y=mx+b ^ ^ slope y-intercept okay so m= slope and b = y-intercept y=mx+b , m=7/2 , (-8,-11) <--- these are the points u will use to find the y-intercept so u will substitute -8 by x and -11 by y so it becomes -11=7/2(-8)+b -11-56/2=b , -22/2-56/2=b , -78/2=b , -39=b okay so we found the y-intercept which is -39 so the equation is y=7/2x-39 i hope u understand...!!!! :D :D

OpenStudy (anonymous):

yes, I do thank you.